### v1

# AutoGuide: Automated Generation and Selection of State-Aware Guidelines for Large Language Model Agents

Abstract
The primary limitation of large language models (LLMs) is their restricted understanding of the world. This poses significant difficulties for LLM-based agents, particularly in domains where pre-trained LLMs lack sufficient knowledge. In this paper, we introduce a novel framework, called AutoGuide, that bridges the knowledge gap in pre-trained LLMs by leveraging implicit knowledge in offline experiences. Specifically, AutoGuide effectively extracts knowledge embedded in offline data by extracting a set of state-aware guidelines. Importantly, each state-aware guideline is expressed in concise natural language and follows a conditional structure, clearly describing the state where it is applicable. As such, the resulting guidelines enable a principled way to provide helpful knowledge pertinent to an agent’s current decision-making process. We show that our approach outperforms competitive LLM-based baselines by a large margin in sequential decision-making benchmarks.
[…]
[0]leftmargin=*
1 Introduction
Recent advances in large language models (LLMs) have empowered AI agents to address various sequential decision-making tasks and applications (Wang et al., 2023; Xi et al., 2023). The foundation of these successes involves the planning and reasoning capabilities of pre-trained LLMs, enabling agents to execute effective policies (Brohan et al., 2023; Wei et al., 2022). However, the knowledge embedded in pre-trained LLMs is confined to the concepts and information present in a training dataset, which is inherently incomplete (Yin et al., 2023). This limited understanding in pre-trained LLMs poses significant difficulties for agents, particularly in target domains with insufficient prior knowledge. For example, in web navigation, LLM agents generally achieve low success rates due to a lack of information about taking appropriate actions on real-world websites with diverse and dynamic contents (Koh et al., 2024; Deng et al., 2023; Gur et al., 2023; Zhou et al., 2023).
In this paper, we explore data-driven strategies that leverage offline experiences to bridge knowledge gaps between pre-trained LLMs and a downstream domain. As offline experiences implicitly convey valuable knowledge about desirable and undesirable policies in domains, they promise to serve as a useful resource for improving an LLM agent’s decision-making in situations where the pre-trained LLM lacks understanding. Despite this potential benefit, a critical challenge lies in effectively extracting the implicit information embedded in offline data. Unfortunately, the conventional method of directly providing all available experiences to an agent, as in in-context learning, is either impossible or unsuccessful due to constraints, including context length limitations, prompt sensitivity, and difficulty with complex reasoning (Lu et al., 2022; Dong et al., 2022; Min et al., 2022; Kaddour et al., 2023).
To address the challenge of extracting knowledge from offline data, we propose a novel framework, called AutoGuide. Specifically, AutoGuide automatically extracts a comprehensive set of state-aware guidelines from offline experiences. Our method then applies these state-conditional guidelines to improve an LLM agent’s performance by identifying the current state and then incorporating its corresponding guidelines into the prompt during testing (see Figure 1). Notably, we generate state-aware guidelines in concise natural language statements, efficiently compressing knowledge in offline data. Additionally, state-aware guidelines clearly describe the states where each guideline is applicable, so AutoGuide provides an LLM agent with guidelines that are most relevant to its current decision-making process. As a result, AutoGuide achieves the highest success rates compared to competitive baselines in challenging sequential decision-making benchmark environments.
Figure 1: AutoGuide aims to extract the implicit knowledge embedded in offline experiences and help the decision-making process of an LLM agent. Specifically, our method generates a comprehensive set of state-aware guidelines from offline data and explicitly identifies when each guideline is applicable by generating its corresponding state. Our state-aware guidelines enable providing pertinent guidelines at test time by summarizing the current trajectory into the state, leading to correct decision-making compared to baselines without state-aware guidelines.
Our contribution. In summary, we present the following main contributions in this paper:
  • Principled method based on state-aware guidelines (Section 3): We develop two modules to generate state-aware guidelines from offline experiences: the state summarization module for generating a description of a given state, and the guideline extraction module for extracting a desired guideline corresponding to the state. The outcome is a set of domain knowledge in concise natural language that enhances decision making by providing pertinent information.
  • Comprehensive evaluation of AutoGuide (Section 4.2): We show AutoGuide’s capability in extracting helpful state-aware guidelines in various interactive benchmark domains, including navigating realistic web environments. Our results highlight the flexibility of AutoGuide, such that it can be seamlessly incorporated into an existing LLM-based method.
  • Analyses with important perspectives (Section 4.3): We explore various aspects of AutoGuide, such as the significance of determining the applicability of each guideline based on generated states. We also investigate having only successful or unsuccessful trajectories in offline experiences, and demonstrate that AutoGuide can still extract valuable knowledge in these corner cases.
2 Related Work
LLM agents. Language models have recently been shown to possess strong priors for sequential decision-making tasks, which has given rise to LLM-powered agents (Wang et al., 2023; Xi et al., 2023; Zheng et al., 2024; Zeng et al., 2023). Agents need to possess various skills to be effective in practice including planning (Brohan et al., 2023; Huang et al., 2022; Logeswaran et al., 2022), reasoning (Wei et al., 2022; Gao et al., 2023), tool manipulation (Qin et al., 2023; Patil et al., 2023; Parisi et al., 2022; Schick et al., 2023), code generation (Sun et al., 2023; Logeswaran et al., 2022), among others. In this work we focus on building effective agents for web (Yao et al., 2022b; Zhou et al., 2023) and embodied (Shridhar et al., 2021) environments.
Self-reflection from past experiences. An important capability for agents to succeed is the ability to learn from past experiences and update their behavior based on feedback. Self-feedback (Madaan et al., 2023; Kim et al., 2023; Shinn et al., 2023) has emerged as an effective technique where a model inspects its own incorrect predictions, reflects on it to identify what went wrong and attempts to improve its prediction. Our approach shares similar motivations to self-refinement approaches, but unlike these methods which generate reflections on the fly, our approach first attempts to distill a set of guidelines from offline experiences and then retrieves the appropriate guidelines during inference for more accurate action prediction. However, these advances are complementary to our method and can be used in conjunction with our approach (as shown in the experiments).
Leveraging natural language guidance. Natural Language can be a rich source of information for artificial agents to learn to act efficiently. Prior work has explored the notion of learning from human written text manuals which describe various details about the agent’s environment (Branavan et al., 2012; Hanjie et al., 2021; Zhong et al., 2020). Recent work has explored automatically generating such guidance in the form of chain-of-thought reasoning (Wei et al., 2022; Yao et al., 2022b) which emulates a thought process or rationale for an agent’s predictions. In contrast to ReAct, which generates such guidance on the fly during inference, our approach examines successful and unsuccessful trajectories corresponding to a task to generate appropriate guidance and uses these guidelines for predicting better actions. ExpeL (Zhao et al., 2023) proposed a related approach to derive guidelines. In contrast to ExpeL, where all guidelines are provided to an agent as a prompt, our guideline selection process is contextual, where guidelines relevant to the agent’s current state are retrieved and used for prediction. We show that this substantially improves over ExpeL’s non-contextual guideline-based approach.
Figure 2: State-aware guideline generation process based on a pair of success trajectory $\bm{\tau}^{i}_{\text{success}}$ and failure trajectory $\bm{\tau}^{i}_{\text{fail}}$. In this example, the two trajectories start deviating from each other at $t=1$. The state summarization module generates a description of a state at $t=1$ given $\bm{\tau}^{i}_{:t}$, and the guideline extraction module generates the corresponding guideline for that state.
3 AutoGuide: Principled Method Based on State-Aware Guidelines
Our work is motivated by the growing availability of offline experiences that agents or humans naturally accumulate during their interactions with the environment. We focus on leveraging these offline experiences to enhance the decision-making of an LLM agent. A key question in achieving our goal is how to effectively extract implicit knowledge embedded in these experiences. This section details our method, AutoGuide, which extracts a comprehensive set of state-aware guidelines from offline data.
3.1 Problem Statement
The goal of AutoGuide is to distill knowledge from offline data into a helpful natural language format, such that the extracted information improves the decision-making of an agent during testing. Formally, AutoGuide is given offline experiences $\mathcal{D}_{\text{train}}\!=\!(\bm{\tau}^{1},...,\bm{\tau}^{N})$ that consist of $N$ trajectories from the training tasks. Each trajectory $\bm{\tau}^{i}\!:=\!(x_{0},a_{0},r_{0},...,r_{T})$ is a sequence of observations, actions, and rewards following the Markov decision process (Sutton & Barto, 2018). We assume that each trajectory $\bm{\tau}^{i}$ can be categorized as either a success or failure determined by either direct feedback from the environment or the terminal reward (e.g., considered a success if the final reward $r_{T}$ is positive) depending on the domain.
3.2 Constructing Dataset for Guideline Extraction
From an offline experience dataset $\mathcal{D}_{\text{train}}$, we construct a paired offline dataset $\mathcal{D}_{\text{paired}}$ for guideline extraction:
$\begin{split}\mathcal{D}_{\text{paired}}\!&=\!\big{(}(\bm{\tau}^{1}_{\text{success}},\bm{\tau}^{1}_{\text{fail}}),...,(\bm{\tau}^{n}_{\text{success}},\bm{\tau}^{n}_{\text{fail}})\big{)},\end{split}$ (1)
where $\bm{\tau}^{i}_{\text{success}}\!\in\!\mathcal{D}_{\text{train}}$ and $\bm{\tau}^{i}_{\text{fail}}\!\in\!\mathcal{D}_{\text{train}}$ denote a success and failure trajectory in $\mathcal{D}_{\text{train}}$ of the same task, respectively. In practice, we can easily collect these pairs by applying advanced LLM-based planning approaches (Yao et al., 2022b; Shinn et al., 2023) or taking advantage of human demonstration datasets (Zhou et al., 2023; Deng et al., 2023). However, in rare cases, it can be challenging to obtain paired data if the offline data $\mathcal{D}_{\text{train}}$ contains only one type (either success or failure) of the trajectories. We discuss how our algorithm can be adjusted to handle such a biased dataset and its impact on the performance in Section 4.3.
3.3 Extraction of State-Aware Guidelines
AutoGuide generates state-aware guidelines by utilizing pairs of success and failure trajectories from offline data. Each state-aware guideline is expressed in concise natural language and follows a conditional structure, clearly describing the state in which this guideline is applicable. Intuitively, contrasting actions taken between $\bm{\tau}^{i}_{\text{success}}$ and $\bm{\tau}^{i}_{\text{fail}}$ provides important information about when and which actions are effective or ineffective. Building on this insight, we develop two modules for deriving state-aware guidelines in a conditional structure (see Figure 2):
Algorithm 1 Collection of state-aware guidelines
  Input: Paired offline dataset $\mathcal{D}_{\text{pair}}$, state summarization module $\mathcal{M}_{\mathrm{state}}$, guideline extraction module $\mathcal{M}_{\mathrm{guideline}}$
  Initialize state-aware guideline dictionary $\mathcal{G}$
  for Each pair ($\bm{\tau}^{i}_{\text{success}},\bm{\tau}^{i}_{\text{fail}})\in\mathcal{D}_{\text{pair}}$  do
     # Summarize a trajectory into a state
     Identify target timestep $t$ from $\bm{\tau}^{i}_{\text{success}}$ and $\bm{\tau}^{i}_{\text{fail}}$
     $\texttt{state}\leftarrow\mathcal{M}_{\mathrm{state}}(\bm{\tau}^{i}_{:t})$
     # Check if the current state matches any existing states
     if $\texttt{state}\notin\mathcal{G}$ then
        $\mathcal{G}[\texttt{state}]=\{\}$
     end if
     # Generate the corresponding state-aware guideline
     $\texttt{guideline}\leftarrow\mathcal{M}_{\text{guideline}}(\bm{\tau}^{i}_{\text{success}},\bm{\tau}^{i}_{\text{fail}},\texttt{state})$
     $\mathcal{G}[\texttt{state}]\leftarrow\mathcal{G}[\texttt{state}]\cup\{\texttt{guideline}\}$
  end for
  Return State-aware guideline dictionary $\mathcal{G}$
State summarization module. The objective of this module is to generate a concise description of the state of a trajectory in natural language. Specifically, given a pair of $\bm{\tau}^{i}_{\text{success}}$ and $\bm{\tau}^{i}_{\text{fail}}$ from the same task, we first identify the target timestep $t$ at which these two trajectories start deviating from each other with different actions. Then, we construct a target trajectory $\bm{\tau}^{i}_{:t}\!:=\!(x_{0},a_{0},...,x_{t})$, which comprises a sequence up to $t$, either from a success or failure trajectory and prompt an LLM to summarize it:
$\begin{split}\text{{state}}\!\leftarrow\!\mathcal{M}_{\text{state}}(\bm{\tau}^{i}_{:t}),\end{split}$ (2)
where Section B.1 details our prompt template. For example, Figure 2 shows trajectories related to the task $i$ of creating a discussion post about online learning in a relevant subReddit. Given that these two trajectories have different actions at a target timestep of $t\!=\!1$, our state summarization module generates the following state summary from $\bm{\tau}^{i}_{:1}$ in Figure 2: “You are on the List of forums page”.
Guideline extraction module. The goal of this module is to generate a desired guideline corresponding to the state. We extract a useful natural language guideline by examining successful and failure trajectories with respect to the state:
$\begin{split}\text{{guideline}}\!\leftarrow\!\mathcal{M}_{\text{guideline}}(\bm{\tau}^{i}_{\text{success}},\bm{\tau}^{i}_{\text{fail}},\texttt{state}),\end{split}$ (3)
where we refer to Section B.2 for the prompt template we use. As an example, this module generates the following state-aware guideline for the state described in Figure 2: “When you are on the List of forums page, … you can click on the link that exactly matches the forum name …”.
Algorithm 2 Applying state-aware guidelines at test time
  Input: State-aware guideline dictionary $\mathcal{G}$, state summarization module $\mathcal{M}_{\text{state}}$, guideline selection module $\mathcal{M}_{\text{select}}$, LLM agent policy $\pi$
  Initialize test trajectory $\bm{\tau}=\{x_{0}\}$
  for Each timestep $t$ do
     # Summarize a trajectory into a state
     $\texttt{state}\leftarrow\mathcal{M}_{\text{state}}(\bm{\tau})$
     # Check if the current state matches any existing states. If matched, then perform top-$k$ guideline selection
     if $\texttt{state}\in\mathcal{G}$ then
        $\texttt{guidelines}\leftarrow\mathcal{M}_{\text{select}}(\texttt{state},\mathcal{G}[\texttt{state}],\bm{\tau})$
     else
        $\texttt{guidelines}\leftarrow\varnothing$
     end if
      # Action selection based on guidelines
     $a_{t}\sim\pi(\bm{\tau},\texttt{state},\texttt{guidelines})$
     Execute action $a_{t}$ and observe $x_{t+1}$
     Update trajectory $\bm{\tau}\leftarrow\bm{\tau}\cup\{\texttt{state},a_{t},x_{t+1}\}$
  end for
Construction of state-aware guidelines. We collect state-aware guidelines $\mathcal{G}$ by iterating through available pairs in the paired offline data and organize the guidelines in a dictionary format, using the state as the key and the corresponding guidelines as the value (see Algorithm 1). In particular, we observe that the state summarization module occasionally produces states that share the same context but are expressed slightly differently. To minimize redundancy, we employ an LLM to determine if the current state corresponds to any previously identified state. If a match is found, we reuse the existing state; otherwise, we introduce the new state into our dictionary. The specific prompt template for this state-matching procedure is outlined in Section B.3.
Figure 3: Sequential decision-making benchmark domains considered in our work: ALFWorld (Shridhar et al., 2021), WebShop (Yao et al., 2022a), WebArena (Zhou et al., 2023). Graphic credit: Shridhar et al. (2020); Yao et al. (2022a); Zhou et al. (2023).
3.4 Applying State-Aware Guidelines at Test Time
After extracting a set of state-aware guidelines $\mathcal{G}$ from offline experiences, our method employs these guidelines to enhance the decision-making of an LLM agent during testing. At each timestep, AutoGuide summarizes the current test trajectory $\bm{\tau}$ into a state using our state summarization module $\mathcal{M}_{\text{state}}$. Then, we employ the LLM designed for identifying a matching state in Section 3.3 to find a match and retrieves its corresponding guidelines. If there are more than $k$ guidelines related to the current state, we apply the guideline selection module to choose top-$k$ guidelines:
$\begin{split}\mathcal{M}_{\text{select}}(\texttt{state},\mathcal{G}[\texttt{state}],\bm{\tau}),\end{split}$ (4)
where Section B.4 details the prompt template for this selection procedure. Subsequently, AutoGuide incorporates both the state and guidelines into the agent’s prompt. Finally, the agent selects an action by considering the provided state and guidelines (see Figure 1 for an example). This process iterates until the end of the test trajectory (see Algorithm 2 for pseudocode).
Key benefits of AutoGuide. First, the extraction of state-aware guidelines in AutoGuide offers the inherent benefit of providing relevant guidelines for the state of interest. This capability is important since neglecting the specific context in which a guideline applies can confuse the agent’s decision-making process. The second key benefit is the generation of concise natural language guidelines, which can be seamlessly incorporated into any prompt-based LLM agent. Lastly, AutoGuide generates guidelines at the individual state level rather than at the trajectory level. Given that a single incorrect action can lead to a complete failure, it is essential to provide detailed assistance in each action selection process. With these advantages, we demonstrate in the next section that our approach significantly enhances the performance of LLM agents.
4 Evaluation
We demonstrate the efficacy of AutoGuide by conducting experiments on a diverse suite of sequential decision-making benchmark domains. We also perform important analyses about AutoGuide. Additional experimental details and hyperparameters can be found in Appendix A.
4.1 Evaluation Setup
4.1.1 Decision-Making Benchmark domains
We consider the following interactive text-based sequential decision-making benchmarks to study different aspects of AutoGuide (see Figure 3):
  • ALFWorld (Shridhar et al., 2021): In this embodied benchmark, an LLM agent interacts with an environment to carry out household tasks, such as placing a pan on the dining table. Observations and actions are expressed in natural language statements, and the agent must navigate through the space and manipulate objects to successfully complete the tasks.
  • WebShop (Yao et al., 2022a): This interactive web environment simulates the task of online shopping on an e-commerce website. The agent’s goal is to understand a text instruction and buy a product that meets specified criteria. This involves querying the website’s search engine, understanding the descriptions and details of each item, and selecting necessary options.
  • WebArena (Zhou et al., 2023): This web-based benchmark introduces realistic environments by replicating the functionality and data found in popular web domains (e.g., Gitlab, Reddit, Wikipedia). Compared to WebShop, WebArena presents more challenges and difficulties for an LLM agent due to its large observation and action space, along with tasks that involve longer planning horizons. We focus on the Reddit domain for the WebArena experiments.
Table 1: Test reward and success rate on WebShop. Agent model for all methods is basedon GPT-3.5-turbo. Reflexion is done by GPT-4-turbo for at most 3 trials.
Table 2: Test result on ALFWorld.
[…]
Algorithm WebArena SR
ReAct $8.0$%
ExpeL $21.8$%
AutoGuide $43.7$%
Table 1: Test reward and success rate on WebShop. Agent model for all methods is basedon GPT-3.5-turbo. Reflexion is done by GPT-4-turbo for at most 3 trials.
Table 2: Test result on ALFWorld.
Table 3: Test result on WebArena.
4.1.2 Baselines
We compare AutoGuide against the following baseline approaches to fully show the effect of state-aware guidelines (refer to Appendix A for more details):
  • ReAct (Yao et al., 2022b): This LLM-based planning method integrates reasoning and acting to address sequential decision-making tasks. However, it does not leverage offline experiences and thus suffers from the limited understanding of pre-trained LLMs in downstream domains.
  • ExpeL (Zhao et al., 2023): This method also extracts natural language knowledge from offline data. However, it fails to consider the applicability of guidelines and does not generate state-aware guidelines. Instead, it provides all guidelines to an LLM agent without filtering out irrelevant ones based on the current state. ExpeL has two contributions, the guideline generation and in-context example selection module. Because the latter is orthogonal to our analysis and can be seamlessly combined with our method, we consider ExpeL with guidelines in our experiments.
  • Reflexion (Shinn et al., 2023): This approach converts environmental feedback into text statements to assist an LLM agent (e.g., ReAct) in the next trial. The baseline generates valuable feedback about solving a specific test task. We demonstrate how state-aware guidelines derived by AutoGuide can be combined with the feedback.
4.1.3 Implementation
We collect offline experiences either by running ReAct and Reflexion, or incorporating human demonstrations. We use ReAct with GPT-3.5-turbo as our base LLM agent for WebShop and ALFWorld and GPT-4-turbo for WebArena. For each benchmark, we apply the same GPT version for action generation, state summarization, and guideline selection. We extract state-aware guidelines from the training set with GPT-4-turbo and evaluate their effectiveness by applying them to the test set with non-overlapping tasks. We refer to Appendix A for more details.
Figure 4: Trajectories of ReAct, ExpeL, and AutoGuide from the same test task. ReAct (Left) chose the wrong item, consequently failing the task in the end. ExpeL (middle) was confused by guidelines that were irrelevant to its state, leading to incorrect reasoning and actions. AutoGuide (right) can select useful guidelines relevant to the agent’s state, enabling the agent to efficiently accomplish the task.
4.2 Main Results
Q1. How effective is AutoGuide compared to the competitive baseline methods without state-aware guidelines?
To answer this question, we run ReAct, ExpeL, and AutoGuide on WebShop and ALFWorld benchmarks. The performance on the test datasets is presented in Tables 3 and 3. There are three notable observations:
  1. 1.
    Effectiveness of state-aware guidelines. Our approach surpasses baseline performance in both benchmarks, achieving the highest test rewards and success rates in Tables 3 and 3. These results highlight the effectiveness of employing state-aware guidelines in language-based decision-making domains. To further examine the action selection differences among ReAct, ExpeL, and our method, we present their trajectories in Figure 4. We observed that ReAct makes common mistakes such as trying to take soapbar that is not visible, or taking a soapbottle instead of soapbar due to their similar names. Both ExpeL and AutoGuide improve on this by extracting guidelines from similar mistakes in the offline experience. However, ExpeL often erroneously applies incorrect guidelines due to the availability of all guidelines at each timestep. In Figure 4, ExpeL mistakenly attends to the second guideline “ensure to specify the item’s number and location…”, leading to wrong reasoning and action. AutoGuide presents relevant guidelines at necessary moments, enabling accurate task completion by avoiding the mistakes seen in ExpeL and ReAct.
  2. 2.
    Limitation of in-context learning. In Table 3, while increasing the number of in-context examples (i.e., #-shot) for ReAct gradually improves performance, there is a plateau at a certain number of shots. Additionally, ReAct with more than 6 shots often exceeds the token limit of GPT-3.5-turbo. These results indicate that directly inputting raw trajectories into ReAct for in-context learning is not an effective way to fully leverage offline data. In contrast, AutoGuide extracts knowledge from entire training trajectories by summarizing them into concise state-aware guidelines, making them easy to integrate with prompt-based agents.
  3. 3.
    Importance of providing pertinent knowledge. ExpeL approach helps ReAct by extracting knowledge from offline experiences, but its impact is not as significant as AutoGuide. Recall that for ExpeL, the guidelines are neither generated for specific states at training time, nor selected to only provide state-aware guidelines at test time. As a result, irrelevant guidelines can be introduced to an agent, potentially causing confusion for the agent. Consequently, the result highlights the significance of providing relevant guidelines conditioned on states for LLM agents.
Q2. Does AutoGuide scale to more challenging and complex environments?
In this question, we highlight the scalability of AutoGuide in addressing challenging and complex domains. Specifically, we conduct experiments on WebArena-Reddit, which features more diverse tasks on realistic and complex websites requiring longer action sequences. It has the larger observation space and more complex action space (e.g., scrolling). Table 3 presents the results, where AutoGuide achieves the highest success rate with a significant margin when compared to ReAct and ExpeL. We observe that ReAct scores low task success rate (8.0%) in WebArena due to the complex observation and action spaces and longer task horizon. In ExpeL, the issue of presenting all guidelines to an agent is exacerbated in the WebArena compared to simpler environments like ALFWorld and WebShop. WebArena’s wide variety of tasks across different domains requires a larger number of guidelines to cover the knowledge needed for all tasks and domains. This results in either an overload of irrelevant guidelines that could mislead the agent or a lack of crucial information when the number of guidelines is limited, as suggested in ExpeL (Zhao et al., 2023). In contrast, AutoGuide achieves a more significant performance enhancement (43.7%) compared to ExpeL (21.8%) by efficiently providing pertinent guidelines for each state and minimizing the burden on context capacity. Refer to Figure 15 for a list of example states and guidelines.
Q3. How does AutoGuide perform when combined with test-time self-feedback approaches like Reflexion?
Our state-aware guidelines effectively provide inter-task knowledge by considering multiple tasks in offline data. Meanwhile, self-feedback methods (e.g., Reflexion) offer intra-task knowledge based on environmental feedback during test time. In this question, we explore the effectiveness of integrating both inter-task and intra-task information. The results presented in Table 3 demonstrate that the combination of AutoGuide with Reflexion achieves the highest performance in the WebShop benchmark. Hence, we find that our state-aware guidelines positively complement the intra-task knowledge of Reflexion. Another observation from Table 3 is that, while ExpeL + Reflexion outperforms ExpeL alone, this combination is not as effective as other approaches. This limitation may stem from ExpeL introducing irrelevant knowledge, potentially leading to conflicts with Reflexion’s feedback and having an adverse impact on the decision-making process.
Figure 5: Extracting state-aware guidelines when offline experiences have either success or failure trajectories. With minor modifications, AutoGuide can extract useful knowledge from offline data and outperform the ReAct baseline.
4.3 Analysis of AutoGuide
Q4. Can AutoGuide extract state-aware guidelines when offline data has either success or failure trajectories?
We explore how AutoGuide can be adapted to address extreme types of offline data, comprising either successful or failed trajectories. Intuitively, successful trajectories exhibit shared patterns of successful behaviors, whereas unsuccessful trajectories may fail for a variety of reasons. Based on this insight, when offline experiences contain exclusively successful instances, we sample multiple trajectories and prompt GPT-4-turbo to identify states exhibiting shared common behaviors across them. Then, for each state, AutoGuide applies the state summarization and guideline extraction modules to construct state-aware guidelines. In the case of failures only, we leverage failed trajectories and a one-shot example from ReAct as a successful demonstration to generate state-aware guidelines based on our modules.
The results are shown in Figure 5. Interestingly, even without having pairs of success and failure trajectories for the same task, the different variants of AutoGuide can extract valuable knowledge from offline data, leading to improvements in test-time performance compared to ReAct, which does not leverage any offline data. As such, our method is flexible and robust to the type of available offline data. However, in general cases, we expect to have both successful and failed trajectories, which provide the most direct insights for AutoGuide to extract useful guidelines.
Q5. How does each component of AutoGuide contribute to the final results?
We evaluate the impact of different components within AutoGuide on its performance in WebShop, as detailed in Table 4. We examine two variants: ReAct+SS and ReAct+GES. The ReAct+SS, which incorporates state summaries into observations without guidelines, shows improvement over ReAct. This suggests that state summaries enhance decision-making by verifying the current state before action selection. ReAct+GES, which generates guidelines from trajectories without state summaries and employs GPT-3.5-turbo for guideline selection, also enhances performance but is less effective than the full AutoGuide. This indicates that choosing relevant guideline based on the trajectory alone is more challenging than using state summary. Therefore, integrating both state summaries and guidelines is crucial for maximizing the benefits of AutoGuide.
[…]
Algorithm SS GES WebShop SR
ReAct 30%
ReAct + SS 36%
ReAct + GES 37%
AutoGuide 46%
Table 4: Ablation study of AutoGuide, analyzing each module’s contribution in the WebShop benchmark. SS denotes our state summarization module, and GES denotes the guideline extraction and selection modules.
5 Conclusion
In this paper, we presented AutoGuide, an effective framework for extracting and exploiting important domain knowledge from offline experiences for improving decision making with pre-trained LLMs. Pre-trained LLMs often lack the knowledge required for solving tasks in downstream domains, and to fill in such knowledge gap, we proposed to generate state-aware guidelines that can be incorporated into prompts for LLM agents. As AutoGuide extracts the guidelines by contrasting successful and failed trajectories on the same task, the resulting state-aware guidelines carry critical information for preventing failures in the domains. For inference, it provides the guidelines pertinent to each of the different states that LLM agents encounter, which can make pre-trained LLMs strong decision-making agents in the downstream domains. Empirically, we showed that AutoGuide outperforms strong LLM-based baselines by a large margin and achieves outstanding performance in multiple sequential decision-making benchmarks.
Appendix A Evaluation Details
A.1 ALFWorld (Shridhar et al., 2021)
A.1.1 Environment details
Each task in ALFWorld starts with a description of the specific environment and the goal to achieve. At each timestep, an agent can choose one of the following actions to interact with the objects and receptacles in the environment:
  • go to [recep]
  • take [object] from [recep]
  • put [object] in/on [recep]
  • open/close/use [recep]
  • clean/heat/cool [object] with [recep]
Alternatively, the agent can generate think actions for planning and reflection, which helps with decision-making but does not change the environment itself. After one action is performed, the environment returns an observation that describes view changes.
Following ReAct, we concatenate a list of (observation, action) pairs to show the entire trajectory up to the current timestep for LLM agents to generate the next action. We experiment on 134 unseen test tasks with 6 categories of pick_and_place, pick_clean_then_place, pick_heat_then_place, pick_cool_then_place, look_at_obj, and pick_two_obj. For each task, the agent is allowed to take a maximum of 50 actions.
A.1.2 Baseline and Models
For ALFWorld tasks, we follow the same setting as ReAct by providing 2 in-context examples for each of the 6 task categories. The original results of ReAct in their paper are produced based on Text-Davinci-002. However, this GPT version is no longer available, so we apply gpt-3.5-turbo-instruct instead to generate actions. For ExpeL, we directly take the guidelines from their appendix and append them to the ReAct agent at test time.
A.1.3 Implementation details of AutoGuide
We run the first 100 training tasks of ALFWorld to collect $(\bm{\tau}_{\text{success}},\bm{\tau}_{\text{fail}})$ pairs with ReAct+Reflexion and extract state-dependant guidelines on the collected data. For state summarization, we provide 2-shot demonstrations for each of the 6 task categories. The corresponding prompt templates can be found in appendix B. All parameter details are shown in Table 5.
Parameter name Value
Allowed Episode Length 50
n-shots 2
Agent Model gpt-3.5-turbo-instruct
State Summarization Model gpt-3.5-turbo-instruct
Guideline Selection Model gpt-3.5-turbo-instruct
Guideline Extraction Model gpt-4-1106-preview
top-k guideline selection 2
Table 5: Experiment hyperparameters on ALFWorld. The maximum allowed episode length and n-shots follow the same setup in ReAct.
A.2 WebShop (Yao et al., 2022a)
A.2.1 Environment details
WebShop provides an e-commerce environment, where the objective is to find and buy the product that matches the task-specific Instruction. The agent can select one of the following actions to perform:
  • search[query]
  • click[button]
Following ReAct, the agent can generate think actions to do planning or reflection. After buying a product, the environment returns a reward showing how well the bought product matches the target one in type, price, buying options, and attributes. The reward is calculated by:
$r\!=\!r_{type}\cdot\frac{|U_{att}\cap Y_{att}|\!+\!|U_{opt}\cap Y_{opt}|\!+\!\mathbbm{1}[y_{price}\leq u_{price}]}{|U_{att}|\!+\!|U_{opt}|\!+\!1}$
where y is the bought product and u is the target product. Same as ALFWorld, for WebShop, the agent takes ($\mathrm{obs}_{t}$, $\mathrm{act}_{t}$) pairs for every previous timestep $t$ as input to generate the next action.
A.2.2 Baseline and Models
Following ReAct, experiments are done in a one-shot setting. We apply gpt-3.5-turbo-0613 to generate actions, but when the token number exceeds the token limit (for example, for the n-shot ReAct experiments in Table 3), we use the 16k version of gpt-3.5-turbo-0613 instead. For ExpeL, we could not find how many training tasks the framework used for training. Therefore, we directly apply the guidelines from the appendix of their paper at test time. We only consider ExpeL with guidelines, not ExpeL with in-context example selection in our experiments for a fair comparison. The in-context example selection method is orthogonal to our work and can be easily combined with our method. For Reflexion, as shown in their paper, their 2-shot Reflexion prompt does not work well on WebShop. Therefore, we re-write a zero-shot prompt and apply gpt-4-1106-preview to generate episode-level reflections for all Reflexion experiments. Following Reflexion and ExpeL, the evaluation is done on 100 test tasks. The maximum number of allowed actions for each task is 15. At the same time, each search action shows the top 3 products for the search query. Please refer to Table 6 for more details about the experiments.
A.2.3 Implementation details of AutoGuide
We randomly sample 50 training tasks from the training set of WebShop, on which we run ReAct+Reflexion to collect pairs and generate guidelines. The state summarization prompt is one-shot, which is shown in appendix B. At test time, we ask gpt-3.5-turbo-0613 to select the most relevant top 2 guidelines for each state.
Parameter name Value
Allowed Episode Length 15
# of Search Results 3
n-shots 1
Agent Model gpt-3.5-turbo-0613
State Summarization Model gpt-3.5-turbo-0613
Guideline Selection Model gpt-3.5-turbo-0613
Guideline Extraction Model gpt-4-1106-preview
Reflexion Model gpt-4-1106-preview
top-k guideline selection 2
Table 6: Experiment hyperparameters on WebShop. The maximum allowed episode length, the number of search results per page, and n-shots follow the same setup in ReAct.
A.3 WebArena (Zhou et al., 2023)
A.3.1 Environment details
WebArena provides web-based benchmark environments that closely follow the data and functionality of real-world websites. Unlike other benchmarks like WebShop that provide clean text of website information as observation, WebArena’s webpage content is represented as an accessibility tree, which is a subset of the DOM tree with useful elements of a webpage. For our expeirments, we focus on WebArena Reddit, which simulates the real Reddit websites with users, forums, and posts with abundant text information.
For each task in WebArena, the agent is expected to achieve a task-specific intent. At each timestep, WebArena provides a list of opened tabs, the accessibility tree of the focused webpage, and the URL of the current page as observation. For WebArena, each observation is long. Therefore, following the baseline in WebArena, at each timestep, we only provide the observation of the current timestep to the agent. We additionally provide up to 5 past actions for the agent to understand what it did in the past. The allowed actions in WebArena include the following:
  • goto[url]
  • click[element_id]
  • type[text][1 for enter or 0 for not enter]
  • press[key_combination]
  • scroll [up or down]
  • go_back
The maximum allowed number of actions for a single task is 20. Note that WebArena does not provide training tasks, but the work provides 19 demonstrations for Reddit, each of a different category. Therefore, we set these 19 tasks as the training tasks and then test on the rest 87 tasks.
A.3.2 Baseline and Models
We directly run the two-shot ReAct-style baseline in the official codebase of WebArena using gpt4-preview-1106. For ExpeL, the original paper does not include experiments on WebArena, therefore we try our best to implement our own version and run on the same training tasks as our method.
A.3.3 Implementation details of AutoGuide
Among the 19 human demonstrations provided in WebArena, 17 of them successfully complete the tasks. We directly run ReAct on the successful tasks to collect failure actions and generate guidelines correspondingly. We take advantage of the URL and webpage titles, which are a part of the observation for this benchmark, to help with state summarization.
Parameter name Value
Allowed Episode Length 20
n-shots 2
Agent Model gpt-4-1106-preview
State Summarization Model gpt-4-1106-preview
Guideline Selection Model gpt-4-1106-preview
Guideline Extraction Model gpt-4-1106-preview
top-k guideline selection 2
Table 7: Experiment hyperparameters on WebArena. The number of shots follows the same setup in ReAct.
Appendix B Prompt Templates
B.1 State Summarization
Figure 6: Our prompt template for state summarization (Equation 2 from Section 3.3) in the WebShop domain.
Figure 7: Our prompt template for state summarization (Equation 2 from Section 3.3) in the ALFWorld domain.
Figure 8: Our prompt template for state summarization (Equation 2 from Section 3.3) in the WebArena domain.
We present our prompt templates for state summarization with $\mathcal{M}_{\text{state}}$ (Equation 2 from Section 3.3) for WebShop, ALFWorld, and WebArena in Figures 6, LABEL:, 7, LABEL: and 8, respectively. In ALFWorld, there exist six categories of tasks, and we use state summarization prompting with 2-shot examples for each task, following the practice by Yao et al. (2022b). Figure 7 shows one example for the pick_and_place tasks.
B.2 Guideline Extraction
In Figures 9, LABEL:, 10, LABEL: and 11, we state the prompt templates we use with $\mathcal{M}_{\text{guideline}}$ for extracting guidelines (Equation 3 from Section 3.3) in the WebShop, ALFWorld, and WebArena domains.
Figure 9: Our prompt template for guideline extraction (Equation 3 from Section 3.3) in the WebShop domain.
Figure 10: Our prompt template for guideline extraction (Equation 3 from Section 3.3) in the ALFWorld domain.
Figure 11: Our prompt template for guideline extraction (Equation 3 from Section 3.3) in the WebArena domain.
B.3 State Matching
Figure 12: Our prompt template for state matching (Sections 3.3, LABEL: and 3.4) in all the WebShop, ALFWorld, and WebArena domains.
Figure 12 shows our prompt template for matching the summarized state with one of the existing states if there is any similar state, for the construction of the set of state-aware guidelines (Section 3.3) and the retrieval of relevant guidelines during testing (Section 3.4) in all the three domains: WebShop, ALFWorld, and WebArena.
B.4 Guideline Selection
Figure 13: Our prompt template for selecting most relevant state-aware guidelines during the test time (Equation 4 from Section 3.4) in the WebShop and ALFWorld domains.
Figure 14: Our prompt template for selecting most relevant state-aware guidelines during the test time (Equation 4 from Section 3.4) in the WebArena domain.
For selecting only $k$ most relevant guidelines in case there are more corresponding state-aware guidelines during testing (Equation 4 from Section 3.4), we use the prompt with Figure 13 for WebShop and ALFWorld and Figure 14 for WebArena.
Appendix C Example State-Aware Guidelines
In Figure 15, we show a list of possible states and state-aware guidelines on WebArena.
Figure 15: Example states and corresponding guidelines for WebArena.